# WavLLM: Towards Robust and Adaptive Speech Large Language Model

### v2

@@ 3.2.2 Advanced Multi-Task Training Stage
Given a speech audio clip, we combine different task prompts for this audio segment as well as text-based instruction tuning tasks in the first mixed single-task training stage together as instructions, and then the training target is designed to complete the tasks sequentially and to repeat key parts of each prompt prior to delivering a response. For example, for an utterance in LibriSpeech, ASR, SQA and text-based IT (t-IT) tasks can be combined into multi-task dataset. Please refer to Appendix A.2 for specific examples. In our work, a total of 2.9K hours of various multitask data are used, including ER+t-IT, ASR+t-IT, ST+t-IT, SV+t-IT, SQA+t-IT, ASR+ST, ASR+SQA, ASR+ST+t-IT and ASR+SQA+t-IT combined tasks, which are summarized in Table 1.
@@ 4 Experiments
Table 1: Training data used in the first stage and second stage. For all tasks, the instructions are diverse. “#Hours” refers to the duration of speech data for each task, not the total number of hours of the data source. The targets of SQA tasks are generated using GPT3.5, GPT-4 or LLAMA-2-chat.
Task Description Data Source #Hours
Single -task automatic speech recognition (ASR) LibriSpeech (Panayotov et al., 2015) 960
LibriHeavy medium (Kang et al., 2023) 1800
speech-to-text translation (ST), including English to German (En2De), English to Japanese (En2Ja), and English to Chinese (En2Zh) CoVoST2 (Wang et al., 2020) 440
MuST-C (Di Gangi et al., 2019) 280
speaker verification (SV) VoxCeleb (Nagrani et al., 2017) 1290
emotion recognition (ER)
IEMOCAP Session 1-4 (Busso et al., 2008)
5
speech question answering (SQA), including gender and speaker-related questions, and multi-round QA LibriSpeech 520
AMI (Carletta et al., 2005) 50
Fisher (Cieri et al., 2004) 710
Switchboard (Godfrey et al., 1992) 230
speech question answering (SQA), continue writing tasks LibriSpeech 960
speech question answering (SQA), summary tasks LibriSpeech 410
instruction tuning (IT), including spoken based and text based tasks Alpaca (Taori et al., 2023) 90
Multi -task ER + text based IT IEMOCAP Session 1-4, Alpaca 71
ASR + text based IT LibriSpeech,  Alpaca 274
ST + text based IT CoVoST2, MuST-C, Alpaca 343
SV + text based IT VoxCeleb, Alpaca 243
SQA + text based IT AMI, Fisher,Switchboard, Alpaca 773
ASR + ST LibriSpeech 74
ASR + SQA LibriSpeech 43
ASR + ST + text-based IT CoVoST2, Alpaca 5
ASR + SQA + text-based IT LibriSpeech, Alpaca 1066
4.1 Implementation Details
As mentioned above, the semantic and acoustic speech encoders are the encoder of Whisper-large-v2222https://huggingface.co/openai/whisper-large-v2 and WavLM-base [^3] https://huggingface.co/microsoft/wavlm-base, the backbone LLM is LLaMA-2-chat-7b [^4] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and all of their parameters are frozen. The outputs of both modality adapters have a time stride of 80 ms and a dimension of 2048, and the rank ($R$) of LoRA is set as 32. In the first mixed single-task training stage, the total number of parameters in our model is 7.55 billion, of which 76.6 million are trainable. In the advanced training phase, the bottleneck dimension ($K$) of the prompt adapter is set as 1024. The 4096-dimensional prompt-dependent parameters produced by prompt adapter are element-wise multiplied with the outputs of the LoRA. Our models are trained with the two-stage curriculum-learning method on 32 V100 GPUs using the Adam optimizer, set with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.98$ and batch size equivalent to 30 seconds per GPU, where the first stage consisted of 400K steps and the subsequent stage involved an additional 150K steps. Additionally, we employed a maximum learning rate of $1\times 10^{-4}$, incorporating a warm-up phase for the first 10% of steps. The two-stage training data are presented in data construction part of Section 3.2.
4.2 Evaluation Setup
Corresponding to the training methods, two primary levels of testing tasks were conducted, namely, single-task evaluation and multi-task evaluation. The detailed information of the two types of task evaluations are privided in the Table 2. Single-task evaluation consists of ASR, ST, SV, ER, SQA, and spoken-query-based question answering (SQQA). The key difference between SQQA and SQA is that in SQQA, the audio contains the question, usually about general knowledge, and the model must directly answer the question asked in the speech; whereas, SQA involves speech that ofen takes the form of a conversation or statement, and the model must understand the content of the speech and then integrate with given instructions to answer the question within instructions. In our work, the single-answer multiple-choice questions of English Listening Comprehension examination (Gaokao) in China are used as the zero-shot SQA task, which gives a short dialogue, a question, and three options, the model is required to choose the correct one from the three options. The performance of SQA is not only a measure of the comprehension of the cross-modal speech and textual content, but also serves as an indicator of the model’s generalization capabilities with respect to a diverse array of instructions.
In the multi-task evaluation, two distinct types of tasks are presented, both of which are given a speech audio clip: the tasks that consist of independent instructions (II-Task) and the tasks that feature sequentially progressive instructions, which are also known as CoT tasks. Examples of these two tasks can be found in the Appendix B. For tasks with independent instructions, our focus lies on not only the ability to follow instructions, i.e. instruction following rate (IFR), but also the correct completion of each instruction. Whereas for CoT tasks, our primary concern is the performance of the final instruction, which will be compared to the performance of one-shot non-CoT based instructions. The audio for zero-shot II-task is from the SQA task (Gaokao), the multitasking instruction includes ASR, SQA, ST and the general knowledge question task. The audio for zero-shot CoT task is generated from the Gigaword (Graff et al., 2003) dataset using Microsoft Azure text-to-speech API, and the target German texts are translated from English summaries of Gigaword dataset [^5] Translation directions of ASR+SQA+ST tasks in second advanced training stage are all English to Chinese. by utilizing GPT-4. The CoT task requires the Speech LLM complete the ASR, summary and translation tasks in turn. In contract, the instructions of one-shot non-CoT based tasks require the cross-lingual summarization directly. For open-ended or target-lack test sets, GPT-4 is utilized to score the outputs, including the accuracy of SQQA and II-task, which is conducted three times and then take the average to minimize the randomness introduced by GPT-4. In addition, the IFR of SQQA is scored manually on 10% of the test utterances.
@@ 4.2 Evaluation Setup
Table 2: Single-task and multi-task evaluation benchmarks, including tasks, datasets, and metrics.
@@
ST CoVoST2 En2De BLEU
MUSTC
SV VoxCeleb1 test set Accuracy (%)
ER IEMOCAP Session 5 Accuracy (%)
SQQA WikiQA (Yang et al., 2015) test set Accuracy (%)
SQA MuTual (Cui et al., 2020)
test set of English listening
comprehension exams
Accuracy (%)
Multi-task II-Task In-house, based on MuTual - Accuracy, IFR (%)
@@
Gigaword (Graff et al., 2003)
- R-1, R-2, R-L, BERTScore (Zhang et al., 2019)
@@
4.3 Main Results
Single-task Evaluation
Table 3 shows the main results of our WavLLM on various single-task instructions compared to other open source speech LLMs and baseline system that combines Whisper large-v2 with LLaMA-2-7b-chat in a cascaded fashion. For the ASR task, our chat model achieves state-of-the-art WERs of 2.0% and 4.8% on test-clean and test-other sets of LibriSpeech corpus, surpassing other open-source chat models on the same size (7B), including SALMONN (Tang et al., 2023) and Qwen-Audio-Chat (Chu et al., 2023). Comparable patterns of superior performance are observed in speech-to-text translation, speaker verification, emotion recognition and spoken-query-based question answering tasks.
The SQA task in our paper is the zero-shot English listening comprehension tests in China, which requires the model to be able to accurately understand and follow the prompt, as well as comprehend the content of speech. Two criteria are evaluated: the first requires that the correct choice must be explicitly given to be considered correct, while the second accepts answers that are semantically equivalent to the correct option, and is scored by utilizing the GPT-4 (The score instruction can be found in Appendix C.1). As shown in column “SQA” of Table 3, the first number is the accuracy based on the given option, and the second number is the accuracy scored by GPT-4. The larger the both accuracy, the better the model’s comprehension and generalization capacity, while the smaller the difference between the both accuracy, the better the model’s ability to follow instructions. From the results, we can observe that our WavLLM model consistently surpasses the cascaded Whisper + LLaMA baseline, and other open source speech large language models. Additionally, our WavLLM model supports multiple dialogue scenario, with a representative instance detailed in Appendix E.
@@ 4.3 Main Results
Multi-task Evaluation
Table 4 shows the results of our WavLLM model on zero-shot independent instructions tasks (II-task) and CoT-task instructions compared to other open source speech large language models. Despite the optimization of SALMONN through activation tuning, and the fact that Qwen-Audio-Chat conducts fine-tuning only on audio encoder without impacting LLM by LoRA weights, their performance in following multitasking instructions remains significantly suboptimal. Our final chat model demonstrates a markedly higher instruction-following rate for the II-Task compared to open-source models like SALMONN and Qwen-Audio-Chat (92.50% vs. 24.25%-57.75%), which highlights the necessity and effectiveness of our advanced multi-task training stage with prompt adapter. From the accuracy based on the GPT-4, which focuses not only on whether instructions are followed, but also on whether they are completed correctly, similar trend can be observed (62.44% vs. 19.58%-37.99%). The score instruction can be found the Appendix C.2.
When the model possesses the capability to handle multi-task instructions, we aspire to enhance its performance through the employment of Chain of Thought (CoT) methodology. Specifically, the CoT based prompt is excepted to give a better performance than one-shot non-CoT based instructions which require the cross-lingual summarization directly. We list the examples of these two types of prompts in the Appendix B.2. From the results in Table 4, we can draw two conlcusions: 1) Our WavLLM model produces the best performance on the CoT-task instructions; 2) Compared with the performance given one-shot non-CoT instructions, our model produces consistent performance improvements on all metrics.
Table 4: Multi-task instruction performance of our WavLLM model compared to other open-source speech large language models.
@@ 4.4 Analysis
Table H and 6 show the results of our models after first mixed single-task training stage and second advanced multi-task training stage. As discussed earlier, the model with only mixed training is incapable of generalizing to multi-task instructions, and the poor performance of II-tasks confirms this. After advanced training, our model produces comparable or even better performance on single-task prompts compared to the first-stage model. For zero-shot II-tasks, significant enhancement of generalization ability is obtained after advanced training, as evidenced not only by the increased adherence to instructions (92.50% vs. 26.25%) but also by the higher accuracy of each executed instruction (62.44% vs. 22.92%). For cross-lingual summary tasks using CoT based instructions, our model after the second advanced multi-task training stage consistently outperforms the first stage model. In addition, we found that the first stage model mainly accomplished the ST task and did not perform the summarization task. To better demonstrate the effectiveness of the second stage, we crafted a long story-based CoT task by GPT-4 where the audio contains a 100-word story, and the target is a 20-word summary in German. In this task, if the model solely focuses on completing the translation, there will be a noticeable discrepancy in length between its output and the desired target. From the results of this task in Table 6, the second advanced multi-task training stage model significantly outperforms the first stage model. Similar results are obtained when utilizing SALMONN-7B to conduct story-based CoT task instructions, which produces 10.6, 1.3, 7.8 and 63.90 on R-1, R-2, R-L and BERTScore respectively.
Table 5: Performance of model with or without advanced training on single-task instructions. mixed training means the first mixed single-task training stage, and advanced training means the second advanced multi-task training stage.
@@
test-clean test-other CoVoST2 MUSTC
WER BLEU Acc. Acc. Acc. Acc.
mixed training 2.0 4.8 23.9 21.9 0.91 0.72 0.55 67.30%
+ advanced training 2.0 4.8 23.6 21.7 0.91 0.72 0.57 67.55%
Table 6: Performance of model with or without advanced training on multi-task instructions. II-tasks and CoT tasks refer to independent instruction and chain-of-thought tasks, as detailed in section 4.2.
       Models II-tasks CoT (ASR+SUMM+En2De, gigaword) CoT (ASR+SUMM+En2De, story)
Acc. IFR R-1 R-2 R-L BERTScore R-1 R-2 R-L BERTScore
mixed training 22.92 26.25 14.7 3.3 13.2 69.71 18.0 2.9 13.7 68.61
+ advanced training 62.44 92.50 16.5 4.1 14.7 70.60 24.5 4.8 19.0 72.52
The Effect of Prompt Adapter
Despite the fact that data-level curricular learning benefits the performance on complex cross-modal tasks, we have observed that the method of sharing LoRA parameters between single-task and multi-task instructions can diminish the performance on single-task prompts. Concurrently, we believe that this approach may also constrain the performance on multi-task instructions. A prompt-aware LoRA weight adapter (prompt adapter) is proposed to address this issue. Comparative experiments are conducted to analyze the effect of prompt adapter during the second advanced multi-task training stage. Additionally, we build a one-stage model trained by combining all data, including both single-task and multi-task data. [^6] Due to the computing resource constraints, only a portion of the single-task dataset are utilized during the second advanced multi-task training stage in this section..
From the results of Table 7 and 8, the following conclusions can be drawn. Firstly, the results of two-stage model without a prompt adapter against one-stage model further demonstrate that the two-stage curriculum learning based training is effective as evidenced by 1) the comparable performance of single-task instructions; 2) consistent performance improvements on zero-shot II-task and CoT-task prompts. Secondly, incorporating the proposed prompt adapter consistently outperforms the baseline two-stage model without such module on all single-task and multi-task instructions.
Table 7: Performance of the model across training stages with and without a prompt adapter on single-task instructions. one-stage denotes the model is trained by utilizing all single-task and multi-task data simultaneously. two-stage (LoRA) stands for two-stage training method with only LoRA technique.
       Models ASR ST (En2De) SV ER SQQA SQA
test-clean test-other CoVoST2 MUSTC
@@
Table 8: Performance of the model across training stages with and without a prompt adapter on multi-task instructions.
       Models II-tasks CoT (ASR+SUMM+En2De, gigaword)
Acc. IFR R-1 R-2 R-L BERTScore
@@
WavLM model has been widely used for speech processing as a foundation model, especially for speaker information extraction. Table 9 shows the single-task instruction performance on models with or without WavLM encoder after the first mixed single-task training stage. Incorporating WavLM-base encoder not only brings performance improvements to speaker verification task but also enhances other tasks such as ASR (relative WER reductions of 13.04% and 11.11% on test-clean and test-other) and ST tasks.
Table 9: Single-task instruction performance of models with or without WavLM encoder after the mixed training.
@@
test-clean test-other CoVoST2 MUSTC
@@
In this subsection, the robustness of the speech LLMs is evaluated by comparing the performance between the seen/single and the unseen/diverse prompts on our WavLLM model and SALMONN model [^7] For our WavLLM model, the training prompts are used as the seen prompts, while for SALMONN, the prompts in Fig.4 and 5 in the paper (Tang et al., 2023) are used as the single prompt. Various prompts generated by GPT-4 are used as unseen prompts for our model and diverse prompts for SALMONN.. From the results in Table 10, compared to the SALMONN model, which experienced a decline in performance with unseen or diverse prompts, our WavLLM model does not exhibit any performance degradation with unseen prompts on ASR tasks and even produces performance improvement on the ST task, demonstrating our model’s powerful robustness.
Table 10: Performance of models using seen (or single) or unseen (or diverse) prompts on WavLLM and SALMONN.
Models ASR ST-CoVoST2
WER BLEU
@@
seen/single unseen/diverse seen/single unseen/diverse seen/single unseen/diverse
@@
(a) ASR task with seen and unseen prompts
(b) single tasks with monotonous prompts
(c) single and ASR+ST tasks with monotonous prompts
(d) single tasks with various prompts
Figure 2: TSNE visualization of the proposed prompt adapter’s outputs in our final WavLLM model. Each point corresponds to a prompt.
@@
In this subsection, TSNE (Van der Maaten and Hinton, 2008) based visualization of the proposed prompt adapter’s output is given in Figure 2. Several trends can be observed: 1) The overlap between two clusters of the seen ASR prompts and unseen ASR prompts implies the generalization of the proposed prompt adapter; 2) the clear discrimination among single-task prompts suggests that the proposed prompt adapter is capable of discerning various single-task instructions and assigning disparate weights to each; 3) Similar strong discrimination between single-task and multi-task instructions is obtained which validates our motivation; 4) The wide distribution of the SQA task with various prompts, illustrates that the prompt adapter can accommodate diverse prompts.
5 Conclusion and Future Work

### v3

@@ 3.2.2 Advanced Multi-Task Training Stage
Given a speech audio clip, we combine different task prompts for this audio segment as well as text-based instruction tuning tasks together as instructions. The training target is designed to complete the tasks sequentially and to repeat key parts of each prompt prior to delivering a response. For example, for an utterance in LibriSpeech, ASR, SQA and text-based IT (t-IT) tasks can be combined into multi-task dataset. Please refer to Appendix A.2 for specific examples. In our work, a total of 2.9K hours of various multitask data are used, including ER+t-IT, ASR+t-IT, ST+t-IT, SV+t-IT, SQA+t-IT, ASR+ST, ASR+SQA, ASR+ST+t-IT and ASR+SQA+t-IT combined tasks, which are summarized in Appendix F.
@@ 4 Experiments
Please find implementation details in Appendix G.
@@
Table 1: Single-task and multi-task evaluation benchmarks, including tasks, datasets, and metrics. “Acc.” stands for accuracy.
@@
ST CoVoST2 wang2020covost En2De BLEU
MUSTC di2019must
SV VoxCeleb1 nagrani2017voxceleb test set Acc. (%)
ER IEMOCAP busso2008iemocap Session 5 Acc. (%)
SQQA WikiQA (yang-etal-2015-wikiqa) test set Acc. (%)
SQA MuTual (cui2020mutual) test set Acc. (%)
Multi -task II-Task In-house, based on MuTual - Acc., IFR (%)
@@
Gigaword (graff2003english)
- R-1, R-2, R-L, BERTScore
@@
4.1 Evaluation Setup
Corresponding to the training methods, two primary levels of testing tasks were evaluated, namely, single-task and multi-task evaluations. The detailed information of the two types of task evaluations are provided in the Table 1. Single-task evaluation consists of ASR, ST, SV, ER, SQA, and spoken-query-based question answering (SQQA). The main difference between SQQA and SQA is that in SQQA the questions are directly in the audio, whereas in SQA the questions are given by text-based instructions. In our work, the single-answer multiple-choice questions of English Listening Comprehension examination (Gaokao) in China are used as the zero-shot SQA task, which gives a short dialogue, a question, and three options. The model is required to choose the correct one from three options. The performance of SQA is not only a measure of the comprehension of the cross-modal speech and textual content, but also serves as an indicator of the model’s generalization capabilities with respect to a diverse array of instructions.
In the multi-task evaluation, two distinct types of tasks are tested, both of which are given a speech audio clip: the tasks that consist of independent instructions (II-Task) and the tasks that feature sequentially progressive instructions, which are also known as CoT tasks. Examples of these two tasks can be found in the Appendix B. For II-Task, our focus lies on not only the ability to follow instructions, i.e. instruction following rate (IFR) [^3] The IFR is scored manually on 10% of the test utterances., but also the correct completion of each instruction. Whereas for CoT tasks, our primary concern is the performance of the final instruction, which will be compared to the performance of one-shot non-CoT based instructions. The multitasking instruction of zero-shot II-tasks includes ASR, SQA, ST and the general knowledge question task. The audio for zero-shot CoT task is generated from the Gigaword (graff2003english) dataset using Microsoft Azure text-to-speech API, and the target German texts are translated from English summaries of Gigaword dataset [^4] Translation directions of ASR+SQA+ST tasks in second advanced training stage are all English to Chinese. by utilizing GPT-4. The CoT task requires the Speech LLM to complete ASR, summary and translation tasks in turn. In contract, the one-shot non-CoT based instructions require the cross-lingual summarization directly. For open-ended or target-lack test sets, GPT-4 is utilized to score the outputs, including the accuracy of SQQA and II-task, which is conducted three times and then take the average to minimize the randomness from GPT-4.
Table 3: Multi-task instruction performance of our WavLLM model compared to other open-source speech LLMs.
@@ 4.1 Evaluation Setup
4.2 Main Results
We compare the performance of WavLLM with other open source text-instruction (chat) based speech LLMs, including SALMONN (tang2023salmonn) and Qwen-Audio-Chat (chu2023qwen), as well as the baseline system that cascades Whisper large-v2 with LLaMA-2-7b-chat, across various single-task and multi-task instructions.
Single-task Evaluation
As shown in Table 2, for the ASR task, our chat model achieves state-of-the-art WERs of 2.0% and 4.8% on test-clean and test-other sets of LibriSpeech corpus, surpassing other open-source chat models on the same size (7B). Similar superior performance are observed in ST, SV, ER and SQQA tasks.
The SQA task in our paper is the zero-shot English listening comprehension tests. As shown in column “SQA” of Table 2, two types of accuracy are evaluated: a) the correct option must be explicitly given (the first number); b) answers that are semantically equivalent to the correct option is considered correct (the second number), which are scored by GPT-4 (The scoring instruction can be found in Appendix C.1). The larger the both accuracy, the better the model’s comprehension and generalization capacity, while the smaller the difference between the both accuracy, the better the model’s ability to follow instructions. From the results, we can observe that our WavLLM model consistently surpasses the cascaded Whisper + LLaMA baseline, and other open source speech LLMs (67.55% vs. 25.50-59.30%). Additionally, our WavLLM model supports multiple dialogue scenario, with a representative instance detailed in Appendix E.
Multi-task Evaluation
As shown in Table 3, despite the optimization of SALMONN through activation tuning, and the fact that Qwen-Audio-Chat conducts fine-tuning only on audio encoder without impacting LLM by LoRA weights, their performance in following multitasking instructions remains significantly suboptimal. Our final chat model produces a markedly higher instruction-following rate for the II-Task compared to SALMONN and Qwen-Audio-Chat (92.50% vs. 24.25%-57.75%), which suggests the necessity and effectiveness of our advanced multi-task training stage with prompt adapter. From the accuracy based on GPT-4, which further focuses on whether they are completed correctly, similar trend can be observed (62.44% vs. 19.58%-37.99%). The scoring instruction can be found the Appendix C.2.
When the model is able to handle multi-task instructions, we aspire to enhance its performance by Chain of Thought (CoT) methodology. Specifically, the CoT based prompt is excepted to give a better performance than one-shot non-CoT based instructions. We list the examples of these two types of prompts in the Appendix B.2. From the results in Table 3, we can draw two conclusions: 1) Our WavLLM model produces the best performance on the CoT-task instructions; 2) Compared with the performance given one-shot non-CoT instructions, our model produces consistent performance improvements on all metrics.
Table 4: Model performance with/without advanced training on multi-task instructions. mixed training and advanced training stand for the first and training stage. “BS.” refers to BERTScore zhang2019bertscore.
       Models II-tasks CoT (ASR+SUMM+En2De)
gigaword story
Acc. IFR R-1 R-2 R-L BS. R-1 R-2 R-L BS.
mixed training 22.92 26.25 14.7 3.3 13.2 69.71 18.0 2.9 13.7 68.61
+ advanced training 62.44 92.50 16.5 4.1 14.7 70.60 24.5 4.8 19.0 72.52
@@ 4.3 Analysis
Table 4 shows the results of our models after first mixed single-task training stage and second advanced multi-task training stage [^5] The results of single-task instructions can be found in Appendix H. After advanced training, our model produces comparable or even better performance on single-task prompts compared to the first-stage model.. For zero-shot II-tasks, significant enhancement of generalization ability is obtained after advanced training, as evidenced not only by the increased adherence to instructions (92.50% vs. 26.25%) but also by the higher accuracy of each executed instruction (62.44% vs. 22.92%). For cross-lingual summary tasks using CoT based instructions, our advanced multi-task trained model consistently outperforms the first stage model.
In addition, we found that the first stage model mainly accomplished the ST task and did not perform the summarization task. To better demonstrate the effectiveness of the second stage, we crafted a long story-based CoT task by GPT-4 where the audio contains a 100-word story, and the target is a 20-word summary in German. In this task, if the model solely focuses on completing the translation, there will be a noticeable discrepancy in length between its output and the target. From the results of this task in Table 4, the second advanced multi-task training stage model significantly outperforms the first stage model, up to 65.52% relative improvement on R-2. When compared to SALMONN-7B on story-based CoT task instructions, a similar greater enhancements can be obtained (24.5/4.8/19.0/72.52 vs. 10.6/1.3/7.8/63.90 on R-1, R-2, R-L and BERTScore respectively.).
Figure 2: TSNE visualization of the proposed prompt adapter’s outputs. Each point corresponds to a prompt.
The Effect of Prompt Adapter
Table 5: Model performance across training stages with/without a prompt adapter on single-task instructions. one-stage denotes the model is trained by utilizing all single-task and multi-task data simultaneously. two-stage (LoRA) stands for two-stage training method with only LoRA technique. “t-c”, “t-o”, “CoV.” and “MU.” stand for test-clean, test-other, CoVoST2 and MUSTC.
@@
t-c t-o CoV. MU.
@@
Despite the fact that data-level curricular learning benefits the performance on complex cross-modal tasks, using the same LoRA parameters between single-task and multi-task instructions may diminish the performance on both instructions. A prompt-aware LoRA weight adapter (prompt adapter) is proposed to address this issue. Comparative experiments are conducted to analyze the effect of prompt adapter during the second advanced multi-task training stage. Additionally, we build a one-stage model trained by combining all data, including both single-task and multi-task data [^6] Due to the computing resource constraints, only a portion of the single-task dataset are utilized during the second advanced multi-task training stage in this section..
From the results of Table 5 and 6, the following conclusions can be drawn. Firstly, the results of two-stage model without a prompt adapter against one-stage model further demonstrate that the two-stage curriculum learning based training is effective as evidenced by 1) the comparable performance of single-task instructions; 2) consistent performance improvements on zero-shot II-task and CoT-task prompts. Secondly, incorporating the proposed prompt adapter consistently outperforms the baseline two-stage model without such module on all single-task and multi-task instructions.
Table 6: Model performance across training stages with/without a prompt adapter on multi-task prompts.
       Models II-tasks CoT (gigaword)
Acc. IFR R-1 R-2 R-L BS.
@@
WavLM model has been widely used for speech processing as a foundation model, especially for speaker information extraction. Table 7 shows the single-task instruction performance on models with or without WavLM encoder after the first mixed single-task training stage. Incorporating the weighted sum of all layers in WavLM-base encoder not only brings performance improvements to speaker verification task but also enhances other tasks such as ASR (relative WER reductions of 13.04% and 11.11% on test-clean and test-other) and ST tasks.
Table 7: Single-task instruction performance of models w or w/o WavLM encoder after the mixed training.
@@
t-c t-o CoV. MU.
@@
In this subsection, the robustness of the speech LLMs is evaluated by comparing the performance between the seen and the unseen prompts on our WavLLM model and SALMONN model [^7] Various prompts generated by GPT-4 are used as unseen prompts.. From the results in Table 8, compared to the SALMONN model, which experienced a decline in performance with unseen or diverse prompts, our WavLLM model does not exhibit any performance degradation with unseen prompts on ASR tasks and even produces performance improvement on the ST task, demonstrating our model’s powerful robustness.
Table 8: Model performance using seen(se.) or unseen (unse.) prompts on WavLLM and SALMONN.
Models ASR (WER) ST-CoVoST2 (BLEU)
@@
se. unse. se. unse. se. unse.
@@
In this subsection, TSNE (van2008visualizing) based visualization of the proposed prompt adapter’s output is given in Figure 2. Several trends can be observed: 1) The overlap between two clusters of the seen and unseen ASR prompts implies the generalization of the proposed prompt adapter; 2) the clear discrimination among single-task prompts suggests that the proposed prompt adapter is capable of discerning various single-task instructions and assigning disparate weights to each; 3) Similar strong discrimination between single-task and multi-task instructions is obtained which validates our motivation; 4) The wide distribution of the SQA task with various prompts, illustrates that the prompt adapter can accommodate diverse prompts.
5 Conclusion